### v2

# Anomalous charge density wave in altermagnetism

Abstract
Exploring the intricate interplay between magnetism and charge density waves has long been a fundamental pursuit at the forefront of condensed matter research. In this letter, based on symmetry analysis and first-principles calculations, we propose for the first time that anomalous charge density wave can be realized in two-dimensional altermagnetic WO. The anomalous charge density wave is characterized by three key features: (i) Unlike conventional charge density wave, whose stabilization is driven by the opening of a gap near the Fermi level, the anomalous charge density wave is stabilized by the occupied states with energies shifting lower far away from the Fermi level; (ii) the anomalous charge density wave increases the density of states near the Fermi level and then enhances—rather than diminishes—the metallicity of materials; (iii) altermagnetism plays a crucial role in stabilizing anomalous charge density wave. Thus, our work offers a pathway for exploring both the realization and the underlying mechanisms of anomalous charge density waves in magnetic systems.
Introduction. Altermagnetism, a new magnetic phase distinct from both ferromagnetism and antiferromagnetism, has been proposed theoretically and confirmed experimentally [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Altermagnetic materials exhibit momentum-space anisotropic spin splitting without spin–orbit coupling (SOC) and real-space antiparallel spin alignment with zero net total magnetic moment, thereby sharing the advantages of both ferromagnetic and antiferromagnetic materials. Consequently, numerous exotic phenomena can emerge in altermagnetic materials, such as anisotropic spin splitting leading to spin currents [11, 12, 13, 14], giant and tunneling magnetoresistance [15, 16, 17], and piezomagnetic effects [18]. With SOC, due to the break of the time-reversal symmetry, altermagnetic materials can realize the anomalous Hall [19, 20, 21, 22], anomalous magneto-optical [23, 24], anomalous Nernst [25], anomalous thermal Hall [26] and crystal valley Hall [27] effects. On the other hand, the combination of altermagnetism with other matter phases has also attracted widespread interest. The anisotropic spin-splitting altermagnetism without time-reversal symmetry merging with topology gives rise to new topological phases[21, 28, 29, 30, 31, 32, 33]. Coupling time-reversal-broken altermagnetism with superconductivity can lead to nontrivial superconducting states [34, 35]. Altermagnetic ferroelectricity enables electrical control of magnetism [36, 37, 38, 39].
Charge density wave (CDW) is a common quantum phase that accompanies lattice distortion and generally drives a metal-to-insulator transition or a reduction in the density of states (DOS) near the Fermi level (i.e., poorer metallicity) [40, 41, 42, 43, 44], as illustrated in Fig. 1(a). Due to strong electron-phonon coupling, the suppression of the CDW often precedes the emergence of superconductivity [45, 46]. An intriguing possibility arises if, instead of weakening metallicity, the CDW actually enhances metallicity, thereby transforming the traditionally competitive relationship between CDW and superconductivity into a cooperative one. Conventionally, CDW stabilization stems from the opening of a gap near the Fermi level, which lowers the total energy at the cost of reduced metallicity. If the stability of CDW phase is instead driven by the downward shift of locally occupied bands far from the Fermi level, the metallicity of materials may be enhanced, as illustrated in Fig. 1(b). We term such a CDW as an anomalous CDW. In nonmagnetic materials, CDWs typically degrade metallicity; the introduction of magnetism may facilitate the realization of anomalous CDWs. Very recently, the combination of altermagnetism and CDWs has begun to be explored [47, 48, 49]. Thus, whether or not is the anomalous CDW realized in an altermagnetic material?
In this letter, based on symmetry analysis and first-principles calculations, we propose for the first time that anomalous CDW can be realized in two-dimensional altermagentic WO. Moreover, our calculations reveal that altermagnetism plays a crucial role in stabilizing the CDW in monolayer WO. Thus, our work offers a pathway for exploring both the realization and the underlying mechanisms of anomalous charge density waves in magnetic systems.
Results and analysis. Recently, monolayer CrO and MoO have been predicted to be bipolarized Weyl semimetals with altermagnetism, and they will transition to the Chern insulator phase under strain [50, 21, 51, 29]. Monolayer CrO is a perfect bipolarized Weyl semimetal, while the overlap of the valence and conduction bands in monolayer MoO leads to the presence of both electron and hole Fermi surfaces. Therefore, monolayer MoO exhibits better metallic properties than monolayer CrO. If W atoms substitute Mo atoms, monolayer WO should have stronger metallic properties. Additionally, 5 transition metal oxides may lead to structural distortions [52], so monolayer WO provides the possibility for the coexistence of altermagnetism and anomalous CDW.
Similar to the monolayer CrO, the crystal structure of monolayer WO has the space group symmetry, corresponding to the point group. The primitive cell of monolayer WO contains two W atoms and two O atoms, with an optimized lattice constant of 3.696 Å. Moreover, the detailed calculations reveal that the magnetic ground state of monolayer WO remains -wave altermagnetic across a range of Hubbard U, consistent with monolayer CrO (the results and analyses are provided in the Supplementary Material (SM) [53]). Then, we investigate the dynamical stability of monolayer WO based on altermagnetic structure. We calculate the phonon spectrum of WO, as shown in Fig. 2(a). From Fig. 2(a), WO has two phonon modes with imaginary frequencies respectively at the high-symmetry M and points, and the imaginary frequency at the M point is larger than that at the point. We speculate that the imaginary frequencies might be caused by a CDW. To verify this speculation, we first calculate the phonon spectrum of the supercell. Under the supercell, the M point of the primitive cell Brillouin zone (BZ) is folded to the point, as shown in Fig. 2(b). Therefore, the vibration mode with imaginary frequency at the M point of the primitive cell BZ is folded to the point. Then, we plot these two vibration modes with imaginary frequency, as shown in Fig. 2(c) and 2(d), respectively. From Fig. 2(c) and 2(d), these two vibration modes originate from the vibrations of W atoms. The vibration mode with imaginary frequency at the point is the relative vibration of W atoms along the direction, and the vibration mode with imaginary frequency at the M point is the in-plane relative vibration of W atoms. Based on these two vibration modes, we obtain the crystal structure with lattice distortion through structural optimization, as shown in Fig. 2(e).
After lattice distortion, the symmetry of crystal structure changes from the high-symmetry to the low-symmetry - space group symmetry, corresponding to the point group. Interestingly, the magnetic ground state of monolayer WO with lattice distortion remains -wave altermagnetic (Fig. 2(e)) (the results and analyses are provided in the SM [53]). Then, we calculate the phonon spectrum for this distorted supercell with altermagnetism, as shown in Fig. 2(f). From Fig. 2(f), there are no imaginary frequencies in the phonon spectrum. Note: The tiny imaginary frequency at the point is caused by the finite displacement calculation method, and the density-functional perturbation theory based on the Quantum Espresso [54] can completely eliminate the imaginary frequency at point (the results and analyses are provided in the SM [53]). Therefore, the two imaginary frequencies at the M and points in the normal phase of monolayer WO are indeed caused by a CDW. That is to say, monolayer WO is a material in which CDW and altermagnetism coexist. On the other hand, we also calculated the phonon spectrum of nonmagnetic monolayer WO in the CDW phase, as shown in Fig. 2(g). From Fig. 2(g), there are pronounced imaginary frequencies, reflecting that nonmagnetic monolayer WO does not host a CDW. Therefore, altermagnetism plays a crucial role in stabilizing the CDW in monolayer WO.
A natural question is what causes the CDW in altermagnetic WO. This requires us to further investigate the electronic structure of altermagnetic WO. In the subsequent electronic structure calculations, we include a Hubbard U = 2.2 eV to account for the on-site Coulomb repulsion among the 5 orbitals of W atoms according to previous studies [55, 56]. We first calculate the electronic band structure and density of states of altermagnetic WO without SOC in the normal phase, as shown in Fig. 3(a). From Fig. 3(a), monolayer WO is an altermagnetic semimetal with four pairs of crossing points near the Fermi level. Two pairs of crossing points with opposite polarization are protected by the spin symmetry along the X-M direction, and the other two pairs of crossing points with opposite polarization are protected by the spin symmetry along the Y-M direction. Therefore, monolayer WO is a bipolarized Weyl semimetal, similar to monolayer Fe2WTe4 [29]. On the other hand, since the bands near the Fermi level are mainly contributed by the 5 orbitals of W atoms (Fig. 3(a)), monolayer WO likely exhibits strong SOC effect.
Before calculating the band structure with SOC, we need to first determine the direction of the magnetic easy axis of monolayer WO. This is because the direction of the magnetic easy axis determines the magnetic point group symmetry, which in turn has a significant impact on the electronic structure and topological properties of monolayer WO. Our calculations show that the magnetic easy axis of WO is along the direction. So altermagnetic WO breaks the and symmetries under SOC, which in turn leads to the Weyl points near the Fermi level opening gap. Next, we calculate the electronic structure of monolayer WO with SOC, as shown in Fig. 3(b). Comparing Fig. 3(a) and 3(b), monolayer WO indeed exhibits strong SOC effect, especially the two pairs of Weyl points near the X and Y points open a band gap of 266 meV. However, monolayer WO remains a semimetal, with its corresponding Fermi surface shown in Fig. 3(c). From Fig. 3(c), the hole Fermi surface is near the point, while the electron Fermi surfaces are located near the X and Y points. Clearly, the hole Fermi surface and the electron Fermi surface cannot be nested by the (, ) wave vector, but the electron Fermi surfaces near the X and Y points can be nested by the (, ) wave vector. However, calculation of the Lindhard susceptibility function shows that monolayer WO exhibits no CDW instability driven by Fermi surface nesting (the results and analyses are provided in the SM [53]). In view of the strong phonon softening observed in the normal phase, the CDW in monolayer WO likely arises from the synergy of electron-phonon coupling with altermagnetism.
Is the CDW in monolayer WO an anomalous CDW? To verify it, we calculate the electronic structure of the altermagnetic WO without SOC in the CDW phase, as shown in Fig. 3(d). Unlike the normal phase, the bands along the -X and -Y directions in the CDW phase of WO are spin degenerate, while the bands along the -M direction are spin split (Fig. 3(d)). This is because the spin symmetries and protect the spin degeneracy of the bands along the -X and -Y directions in the CDW phase of monolayer WO. On the other hand, since the states near the Fermi level are still contributed by the 5 orbitals of W atoms, we also calculate the electronic band structure of the CDW phase of WO with SOC (Fig. 3(e)). Although the SOC splits the crossing points along the -M direction and the spin degeneracy on the high-symmetry lines, the splitting is not significant. Importantly, the DOS in the CDW phase is 2.32 times that of the normal phase at the Fermi level, transforming monolayer WO from a semimetal into a metal. Moreover, we present the Fermi surface of the CDW phase of WO, which intuitively shows that the CDW phase of WO becomes a good metal (Fig. 3(f)). Therefore, the anomalous CDW can be realized in altermagnetic WO. In addition, we also calculate the electronic structure of nonmagnetic monolayer WO in the CDW phase (the results and analyses are provided in the SM [53]). In the nonmagnetic phase, monolayer WO remains metallic; its bands differ markedly from those in the altermagnetic phase, yet these states near the Fermi level are still dominated by the 5 orbitals of W atoms. Despite these similarities, the CDW is unstable in the nonmagnetic case. This further underscores the crucial role played by altermagnetism in enabling the anomalous CDW.
Another important question is why the CDW in altermagnetic WO is anomalous. To answer this question, we first compare the crystal-field splittings of the 5 orbitals from W atoms before and after the lattice distortion. In the normal-phase monolayer WO, W sits in a rectangular crystal field whose and orbitals lie at higher energy, while , , and are lower (Fig. 4(a)). After the CDW distortion, W sits in a distorted rectangular field that pushes and even higher and pulls , , and to lower energies (Fig. 4(a)). Moreover, the large downward shift stabilizes the CDW phase. Then, we plot band structures with -orbital projection for normal-phase and CDW-phase WO, as shown in Fig. 4(b) and 4(c), respectively. They confirm the higher-lying and and lower-lying , , and levels, and show that after the CDW transition and rise further while , , and down further (Fig. 4(b) and 4(c)). Notably, the bands crossing the Fermi level are dominated by the and orbitals (Fig. 4(b) and 4(c)); their upward shift in the CDW phase enlarges the Fermi surface and turns WO into a metal. Therefore, the origin of the anomalous CDW in monolayer WO is fully consistent with the proposed mechanism illustrated in the schematic (Fig. 1(b)).
According to the results and analysis above, the anomalous CDW in monolayer WO exhibits three distinctive features: (i) Altermagnetism plays a crucial role in stabilizing anomalous CDW; (ii) Unlike conventional CDW, whose stabilization is driven by the opening of a gap near the Fermi level, the anomalous CDW is stabilized by the occupied states with energies shifting lower far away from the Fermi level; and (iii) The second feature increases the density of states near the Fermi level, thereby driving the anomalous semimetal-to-metal transition in monolayer WO induced by CDW. Furthermore, the three distinctive features of the anomalous CDW in monolayer WO are not confined to altermagnetic materials; they may also extend to other magnetic phases, such as collinear antiferromagnetism and unconventional non-collinear antiferromagnetism. Our work therefore offers a promising avenue for exploring anomalous charge density waves in magnetic systems more broadly.
Acknowledgements.
We thank Jun-Wei Liu for valuable discussions. This work was financially supported by the National Natural Science Foundation of China (Grant No.12434009, No.12204533, No.12304165 and No.62476278), the National Key RD Program of China (Grant No. 2024YFA1408601), the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (Grant No. 24XNKJ15). Computational resources have been provided by the Physical Laboratory of High Performance Computing at Renmin University of China. Z.-H. D. and L. W. contributed equally to this work.References
- Šmejkal et al. [2022a] L. Šmejkal, J. Sinova, and T. Jungwirth, Beyond conventional ferromagnetism and antiferromagnetism: A phase with nonrelativistic spin and crystal rotation symmetry, Phys. Rev. X 12, 031042 (2022a).
- Šmejkal et al. [2022b] L. Šmejkal, J. Sinova, and T. Jungwirth, Emerging research landscape of altermagnetism, Phys. Rev. X 12, 040501 (2022b).
- Lee et al. [2024] S. Lee, S. Lee, S. Jung, J. Jung, D. Kim, Y. Lee, B. Seok, J. Kim, B. G. Park, L. Šmejkal, C.-J. Kang, and C. Kim, Broken Kramers Degeneracy in Altermagnetic MnTe, Phys. Rev. Lett. 132, 036702 (2024).
- Krempaskỳ et al. [2024] J. Krempaskỳ, L. Šmejkal, S. D’souza, M. Hajlaoui, G. Springholz, K. Uhlířová, F. Alarab, P. Constantinou, V. Strocov, D. Usanov, et al., Altermagnetic lifting of kramers spin degeneracy, Nature 626, 517 (2024).
- Osumi et al. [2024] T. Osumi, S. Souma, T. Aoyama, K. Yamauchi, A. Honma, K. Nakayama, T. Takahashi, K. Ohgushi, and T. Sato, Observation of a giant band splitting in altermagnetic MnTe, Phys. Rev. B 109, 115102 (2024).
- Reimers et al. [2024] S. Reimers, L. Odenbreit, L. Šmejkal, V. N. Strocov, P. Constantinou, A. B. Hellenes, R. Jaeschke Ubiergo, W. H. Campos, V. K. Bharadwaj, A. Chakraborty, et al., Direct observation of altermagnetic band splitting in CrSb thin films, Nat. Commun. 15, 2116 (2024).
- Ding et al. [2024] J. Ding, Z. Jiang, X. Chen, Z. Tao, Z. Liu, T. Li, J. Liu, J. Sun, J. Cheng, J. Liu, Y. Yang, R. Zhang, L. Deng, W. Jing, Y. Huang, Y. Shi, M. Ye, S. Qiao, Y. Wang, Y. Guo, D. Feng, and D. Shen, Large Band Splitting in -Wave Altermagnet CrSb, Phys. Rev. Lett. 133, 206401 (2024).
- Zhou et al. [2025] Z. Zhou, X. Cheng, M. Hu, R. Chu, H. Bai, L. Han, J. Liu, F. Pan, and C. Song, Manipulation of the altermagnetic order in CrSb via crystal symmetry, Nature 638, 645 (2025).
- Yang et al. [2025] G. Yang, Z. Li, S. Yang, J. Li, H. Zheng, W. Zhu, Z. Pan, Y. Xu, S. Cao, W. Zhao, et al., Three-dimensional mapping of the altermagnetic spin splitting in CrSb, Nat. Commun. 16, 1442 (2025).
- Jiang et al. [2025] B. Jiang, M. Hu, J. Bai, Z. Song, C. Mu, G. Qu, W. Li, W. Zhu, H. Pi, Z. Wei, et al., A metallic room-temperature d-wave altermagnet, Nat. Phys. 21, 754 (2025).
- González-Hernández et al. [2021] R. González-Hernández, L. mejkal, K. Výborný, Y. Yahagi, J. Sinova, T. Jungwirth, and J. elezný, Efficient Electrical Spin Splitter Based on Nonrelativistic Collinear Antiferromagnetism, Phys. Rev. Lett. 126, 127701 (2021).
- Bose et al. [2022] A. Bose, N. J. Schreiber, R. Jain, D.-F. Shao, H. P. Nair, J. Sun, X. S. Zhang, D. A. Muller, E. Y. Tsymbal, D. G. Schlom, and D. C. Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nat. Electron. 5, 267 (2022).
- Bai et al. [2022] H. Bai, L. Han, X. Y. Feng, Y. J. Zhou, R. X. Su, Q. Wang, L. Y. Liao, W. X. Zhu, X. Z. Chen, F. Pan, X. L. Fan, and C. Song, Observation of Spin Splitting Torque in a Collinear Antiferromagnet , Phys. Rev. Lett. 128, 197202 (2022).
- Karube et al. [2022] S. Karube, T. Tanaka, D. Sugawara, N. Kadoguchi, M. Kohda, and J. Nitta, Observation of Spin-Splitter Torque in Collinear Antiferromagnetic , Phys. Rev. Lett. 129, 137201 (2022).
- mejkal et al. [2022] L. mejkal, A. B. Hellenes, R. González-Hernández, J. Sinova, and T. Jungwirth, Giant and Tunneling Magnetoresistance in Unconventional Collinear Antiferromagnets with Nonrelativistic Spin-Momentum Coupling, Phys. Rev. X 12, 011028 (2022).
- Shao et al. [2021] D.-F. Shao, S.-H. Zhang, M. Li, C.-B. Eom, and E. Tsymbal, Spin-neutral currents for spintronics, Nat. Commun. 12, 7061 (2021).
- Zhang et al. [2024] R.-W. Zhang, C. Cui, R. Li, J. Duan, L. Li, Z.-M. Yu, and Y. Yao, Predictable gate-field control of spin in altermagnets with spin-layer coupling, Phys. Rev. Lett. 133, 056401 (2024).
- Ma et al. [2021] H.-Y. Ma, M. Hu, N. Li, J. Liu, W. Yao, J.-F. Jia, and J. Liu, Multifunctional antiferromagnetic materials with giant piezomagnetism and noncollinear spin current, Nat. Commun. 12, 2846 (2021).
- Šmejkal et al. [2020] L. Šmejkal, R. González-Hernández, T. Jungwirth, and J. Sinova, Crystal time-reversal symmetry breaking and spontaneous hall effect in collinear antiferromagnets, Sci. Adv. 6, eaaz8809 (2020).
- Šmejkal et al. [2022] L. Šmejkal, A. H. MacDonald, J. Sinova, S. Nakatsuji, and T. Jungwirth, Anomalous hall antiferromagnets, Nat. Rev. Mater. 7, 482 (2022).
- Guo et al. [2023a] P.-J. Guo, Z.-X. Liu, and Z.-Y. Lu, Quantum anomalous hall effect in collinear antiferromagnetism, npj Comput. Mater. 9, 70 (2023a).
- Hou et al. [2023] X.-Y. Hou, H.-C. Yang, Z.-X. Liu, P.-J. Guo, and Z.-Y. Lu, Large intrinsic anomalous Hall effect in both Nb2FeB2 and Ta2FeB2 with collinear antiferromagnetism, Phys. Rev. B 107, L161109 (2023).
- Zhou et al. [2021] X. Zhou, W. Feng, X. Yang, G.-Y. Guo, and Y. Yao, Crystal chirality magneto-optical effects in collinear antiferromagnets, Phys. Rev. B 104, 024401 (2021).
- Chen et al. [2025] W. Chen, X. Zhou, W.-K. Lou, and K. Chang, Magneto-optical conductivity and circular dichroism in -wave altermagnets, Phys. Rev. B 111, 064428 (2025).
- Han et al. [2025] L. Han, X. Fu, W. He, J. Dai, Y. Zhu, W. Yang, Y. Chen, J. Zhang, W. Zhu, H. Bai, C. Chen, D. Hou, C. Wan, X. Han, C. Song, J. Liu, and F. Pan, Nonvolatile anomalous nernst effect in with a collinear néel vector, Phys. Rev. Appl. 23, 044066 (2025).
- Zhou et al. [2024] X. Zhou, W. Feng, R.-W. Zhang, L. Šmejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal Thermal Transport in Altermagnetic RuO2, Phys. Rev. Lett. 132, 056701 (2024).
- Tan et al. [2025] C.-Y. Tan, Z.-F. Gao, H.-C. Yang, Z.-X. Liu, K. Liu, P.-J. Guo, and Z.-Y. Lu, Crystal valley hall effect, Phys. Rev. B 111, 094411 (2025).
- Antonenko et al. [2025] D. S. Antonenko, R. M. Fernandes, and J. W. F. Venderbos, Mirror chern bands and weyl nodal loops in altermagnets, Phys. Rev. Lett. 134, 096703 (2025).
- Tan et al. [2024] C.-Y. Tan, Z.-F. Gao, H.-C. Yang, K. Liu, P.-J. Guo, and Z.-Y. Lu, Bipolarized weyl semimetals and quantum crystal valley hall effect in two-dimensional altermagnetic materials (2024), arXiv:2406.16603 .
- Gao et al. [2025] Z.-F. Gao, S. Qu, B. Zeng, Y. Liu, J.-R. Wen, H. Sun, P.-J. Guo, and Z.-Y. Lu, Ai-accelerated discovery of altermagnetic materials, Natl. Sci. Rev. 12, nwaf066 (2025).
- Li et al. [2024] Y.-X. Li, Y. Liu, and C.-C. Liu, Creation and manipulation of higher-order topological states by altermagnets, Phys. Rev. B 109, L201109 (2024).
- Feng et al. [2025] P. Feng, C.-Y. Tan, M. Gao, X.-W. Yan, Z.-X. Liu, P.-J. Guo, F. Ma, and Z.-Y. Lu, Type-II quantum spin hall insulator (2025), arXiv:2503.13397 .
- Zhang et al. [2025] R.-W. Zhang, C. Cui, Y. Wang, J. Duan, Z.-M. Yu, and Y. Yao, Quantized Spin-Hall Conductivity in Altermagnet Fe2Te2O with Mirror-Spin Coupling (2025), arXiv:2503.10681 .
- Zhu et al. [2023] D. Zhu, Z.-Y. Zhuang, Z. Wu, and Z. Yan, Topological superconductivity in two-dimensional altermagnetic metals, Phys. Rev. B 108, 184505 (2023).
- Pupim and Scheurer [2025] L. V. Pupim and M. S. Scheurer, Adatom engineering magnetic order in superconductors: Applications to altermagnetic superconductivity, Phys. Rev. Lett. 134, 146001 (2025).
- Duan et al. [2025] X. Duan, J. Zhang, Z. Zhu, Y. Liu, Z. Zhang, I. Žutić, and T. Zhou, Antiferroelectric altermagnets: Antiferroelectricity alters magnets, Phys. Rev. Lett. 134, 106801 (2025).
- Gu et al. [2025] M. Gu, Y. Liu, H. Zhu, K. Yananose, X. Chen, Y. Hu, A. Stroppa, and Q. Liu, Ferroelectric switchable altermagnetism, Phys. Rev. Lett. 134, 106802 (2025).
- Šmejkal [2024] L. Šmejkal, Altermagnetic multiferroics and altermagnetoelectric effect (2024), arXiv:2411.19928 .
- Guo et al. [2023b] P.-J. Guo, Y. Gu, Z.-F. Gao, and Z.-Y. Lu, Altermagnetic ferroelectric LiFe2F6 and spin-triplet excitonic insulator phase (2023b), arXiv:2312.13911 .
- Gruner [2018] G. Gruner, Density waves in solids (CRC press, 2018).
- Grüner [1988] G. Grüner, The dynamics of charge-density waves, Rev. Mod. Phys. 60, 1129 (1988).
- Zhu et al. [2015] X. Zhu, Y. Cao, J. Zhang, E. W. Plummer, and J. Guo, Classification of charge density waves based on their nature, Proc. Natl. Acad. Sci. U.S.A. 112, 2367 (2015).
- Zhu et al. [2017] X. Zhu, J. Guo, J. Zhang, and E. Plummer, Misconceptions associated with the origin of charge density waves, Adv. Phys. X 2, 622 (2017).
- Johannes and Mazin [2008] M. D. Johannes and I. I. Mazin, Fermi surface nesting and the origin of charge density waves in metals, Phys. Rev. B 77, 165135 (2008).
- Chen et al. [2016] C.-W. Chen, J. Choe, and E. Morosan, Charge density waves in strongly correlated electron systems, Rep. Prog. Phys. 79, 084505 (2016).
- Agterberg et al. [2020] D. F. Agterberg, J. S. Davis, S. D. Edkins, E. Fradkin, D. J. Van Harlingen, S. A. Kivelson, P. A. Lee, L. Radzihovsky, J. M. Tranquada, and Y. Wang, The physics of pair-density waves: cuprate superconductors and beyond, Annu. Rev. Condens. Matter Phys. 11, 231 (2020).
- Xu et al. [2025] C. Xu, S. Wu, G.-X. Zhi, G. Cao, J. Dai, C. Cao, X. Wang, and H.-Q. Lin, Altermagnetic ground state in distorted Kagome metal CsCr3Sb5, Nat. Commun. 16, 3114 (2025).
- Regmi et al. [2025] R. B. Regmi, H. Bhandari, B. Thapa, Y. Hao, N. Sharma, J. McKenzie, X. Chen, A. Nayak, M. El Gazzah, B. G. Márkus, et al., Altermagnetism in the layered intercalated transition metal dichalcogenide CoNb4Se8, Nat. Commun. 16, 4399 (2025).
- Candelora et al. [2025] C. Candelora, M. Xu, S. Cheng, A. D. Vita, D. Romanin, C. Bigi, M. B. Petersen, A. LaFleur, M. Calandra, J. Miwa, Y. Hwang, Z. Wang, F. Mazzola, and I. Zeljkovic, Discovery of magnetic-field-tunable density waves in a layered altermagnet (2025), arXiv:2503.03716 .
- Chen et al. [2023] X. Chen, D. Wang, L. Li, and B. Sanyal, Giant spin-splitting and tunable spin-momentum locked transport in room temperature collinear antiferromagnetic semimetallic cro monolayer, Appl. Phys. Lett. 123, 022402 (2023).
- Wu et al. [2023] B. Wu, Y.-L. Song, W.-X. Ji, P.-J. Wang, S.-F. Zhang, and C.-W. Zhang, Quantum anomalous Hall effect in an antiferromagnetic monolayer of MoO, Phys. Rev. B 107, 214419 (2023).
- Khomskii [2014] D. I. Khomskii, Transition metal compounds (Cambridge University Press, 2014).
- [53] See supplemental material at http://link.aps.org/xxx, which includes a detailed description of computational methods as well as supplemental figures.
- Giannozzi et al. [2009] P. Giannozzi, S. Baroni, N. Bonini, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, G. L. Chiarotti, M. Cococcioni, I. Dabo, et al., Quantum espresso: a modular and open-source software project for quantumsimulations of materials, J. Phys. Condens. Matter 21, 395502 (2009).
- Dudarev et al. [1998] S. L. Dudarev, G. A. Botton, S. Y. Savrasov, C. J. Humphreys, and A. P. Sutton, Electron-energy-loss spectra and the structural stability of nickel oxide: An LSDA+U study, Phys. Rev. B 57, 1505 (1998).
- Calderon et al. [2015] C. E. Calderon, J. J. Plata, C. Toher, C. Oses, O. Levy, M. Fornari, A. Natan, M. J. Mehl, G. Hart, M. Buongiorno Nardelli, and S. Curtarolo, The aflow standard for high-throughput materials science calculations, Comput. Mater. 108, 233 (2015).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Togo [2023] A. Togo, First-principles phonon calculations with phonopy and phono3py, J. Phys. Soc. Jpn. 92, 012001 (2023).
- Momma and Izumi [2011] K. Momma and F. Izumi, VESTA3 for three-dimensional visualization of crystal, volumetric and morphology data, J. Appl. Crystallogr. 44, 1272 (2011).
- Ganose et al. [2021] A. M. Ganose, A. Searle, A. Jain, and S. M. Griffin, Ifermi: A python library for fermi surface generation and analysis, J. Open Source Softw. 6, 3089 (2021).
- Hamann [2013] D. R. Hamann, Optimized norm-conserving vanderbilt pseudopotentials, Phys. Rev. B 88, 085117 (2013).
- Baroni et al. [2001] S. Baroni, S. de Gironcoli, A. Dal Corso, and P. Giannozzi, Phonons and related crystal properties from density-functional perturbation theory, Rev. Mod. Phys. 73, 515 (2001).